# Stronger Together: On-Policy Reinforcement Learning for Collaborative LLMs

### v2

@@ 4.2 MAS Training System
Mainstream RL post-training frameworks for LLMs, e.g., TRL (von Werra et al., 2020), VERL (Sheng et al., 2024), AReaL (Fu et al., 2025), and OpenRLHF (Hu et al., 2024) primarily support single-agent RL training, which typically involves: a single agent-environment interaction pattern, a single policy operating on a single data buffer, and a single LLM resource pool. This makes it difficult to (i) train multiple models in on-policy RL, (ii) maintain clean on-policy training data, and (iii) support diverse MAS workflow.
We introduce a novel MAS training system to overcome these challenges and enable AT-GRPO  in Alg. 1. By allocating an independent resource pool to each model, our system is designed to support the concurrent on-policy training of multiple policies. The system, depicted in Fig. 4, consists of the following components:
@@
Each policy is managed within an independent resource pool. Following HybridFlow-style  (Sheng et al., 2025), each pool comprises two workers: a RolloutWorker for inference and an UpdateWorker for optimization. During the rollout phase, all policies interact collectively according to the Alg. 1 and MAS workflow; Once collected, each trajectory is routed to the corresponding UpdateWorker, maintaining an on-policy learning regime for every policy.
@@ 5 Experiments
1. Experimental Setup. We train and evaluate Qwen3 models at 1.7B and 8B in the no-thinking mode (Yang and the Qwen Team, 2025). All runs use a single node with $8\times$ H100 GPUs. The rollout sample size is K=4K{=}4 and the turn horizon is $T=4$. The reward-mixing coefficient is $\alpha{=}1$ without further tuning. Full training details appear in Appendix A.2.1.
@@
We evaluate five variants (all initialized from the same base model): (a) Single Agent (prompt-only): one frozen LLM solves the task end-to-end; (b) Single Agent + GRPO: as (a) but trained with GRPO (Shao et al., 2024); (c) MAS (prompt-only): role-specialized prompting over a frozen, role-sharing backbone; (d) MAS + RL (role-sharing policy): all roles share one policy and pooled trajectories update it jointly; (e) MAS + RL (role-specialized policies): samples are routed by role and each policy is optimized independently (no parameter sharing).
3. Task Setups.
For each task, we compare a multi-agent (MA) workflow to a single-agent (SA) counterpart under the same number of environment workers and a turn budget $K$. Full prompt templates, role instructions, and rule-based rewards are provided in Appendix A.2.2.
Code — MA: Tester builds/refines unit tests; Coder implements/refines code; tests execute every turn; reward is per-case/aggregate pass–fail; termination when alignment is achieved or turns reach $K$. SA: one agent emits code; single-turn termination (no verification loop).
Math — MA: Tool agent issues Python/calculator calls; Reasoner produces the final answer; reward is exact-match or verifier-checked; termination upon success or when turns reach $K$. SA: one agent performs reasoning (with direct tool calls if any) in a single turn; single-turn termination.
Plan/Game — MA: Planner proposes actions; Executor calls tools and returns effects/observations; reward reflects step/terminal goal satisfaction (e.g., Plan-Path, Sudoku, Sokoban); termination when the goal is met or turns reach $K$. SA: one agent outputs a plan (same termination condition).
@@
Sudoku and Sokoban. We evaluate our method on gaming tasks: a $4{\times}4$ Sudoku and a $6{\times}6$ Sokoban. We use instances with an automatic checker, following the symbolic task setup of SymBench (Chen et al., 2025b). To ensure a fair evaluation, we generate distinct training and validation sets using different random seeds and verify there is no overlap.
Plan-Path. We use a $10{\times}10$ grid-based Plan-Path environment. This follows the checker-backed symbolic task setup in CodeSteer’s SymBench (Chen et al., 2025b). To separate training and validation, we generate the two splits with distinct random seeds and verify no duplication.
Code Generation. For training, we adopt size-specific corpora: the 1.7B Qwen model is trained on the APPS training split (introductory-difficulty subset) (Hendrycks et al., 2021), while the 8B model is trained on CodeContests (DeepMind, 2024). For model-generated code, we use the dataset’s golden unit tests to score correctness; for model-generated UT, we use the dataset’s golden reference solutions to compute the reward. For evaluation, we use three widely adopted coding benchmarks spanning interview-style and contest-style settings: APPS (Hendrycks et al., 2021), LiveCodeBench-v6 (White et al., 2024), and CodeContests (DeepMind, 2024).
Mathematical Reasoning. We train on the Polaris-Dataset-53K (An et al., 2025) and evaluate on several standard mathematical reasoning benchmarks. For validation, we use AIME24/AIME25 (Mathematical Association of America & AoPS Community, 2024; 2025) and OlympiadBench (He et al., 2024). All math tasks use verifier-checked numeric scoring.
@@ 5.2 Results and Analysis
This improvement is even more pronounced with the Qwen3 8B model, where MAS + AT-GRPO elevates the success rate from a 14–47% range for the single-agent baseline to 96.0–99.5%. By analyzing the dialogue records between agents, we find this dramatic improvement stems from an emergent collaboration: the tool agent learns to generate correct algorithms (e.g., BFS, $A^{\star}$ search), while the plan agent provides crucial oversight, interpreting execution outcomes and delivering the corrective final action list. On-policy RL training within the MAS enhances inter-agent coordination. Conversely, training agents in isolation results in only limited improvement, as detailed in our ablation study (Sec. 5.3, Tab. 3). Furthermore, on the coding and math benchmarks, our approach yields consistent gains, with absolute gains over the baseline ranging from +2.35 (CodeContests) to +16.30 (APPS) in coding, and from +1.80 (OlympaidBench) to +38.70 (AIME24) in math. We hypothesize two reasons: (1) Base models like Qwen3 have already been extensively trained for these common domains, as noted in their official reports (Yang and the Qwen Team, 2025), potentially leading to performance saturation. (2) The diverse nature of problems within these domains presents a greater challenge for improvement via RL training.
@@ 5.3 Ablation Study
To further investigate the contributions of our core training components, we conducted an ablation study with results summarized in Tab. 3 and Fig. 5. Our analysis yields several observations.
@@
Figure 5: (a) Evolution of standardized rewards for the Tool and Plan agents in the role-specific MAS on Plan-Path with Qwen3 1.7B. Solid curves denote the run-averaged mean rewards; shaded bands show variability across runs. (b) Evolution of the average turns required to solve tasks on coding and math.
@@
6 Conclusion and Discussion
Conclusion. In this paper, we proposed AT-GRPO, an agent- and turn-wise grouped reinforcement learning algorithm tailored for on-policy training in MAS. To support this, we introduced a novel training system capable of managing diverse MAS workflows and performing on-policy updates for multiple policies. Our extensive experiments demonstrate that our method delivers consistent gains across diverse domains. On long-horizon planning tasks, it overcomes a key bottleneck of single-agent RL by boosting accuracy from a 14–47% baseline to 96.0–99.5%. Furthermore, it improves complex reasoning performance, with average gains of 3.87–7.62% on coding and 9.0–17.93% on math tasks. Our analysis reveals that RL training in MAS context reinforces role-specific specialization, with the choice between a shared or specialized policy contingent on the task’s characteristics.
Limitations. While our work demonstrates the effectiveness of on-policy training in MAS, this work focuses exclusively on cooperative tasks. Investigating the adaptability of on-policy RL to mixed-motive or competitive settings remains an important open area. Also, our experiments are confined to text-based environments, a promising future direction is to explore the collaboration between Vision Language Models (VLMs) and LLMs, which is a potential opportunity to unlock new capabilities in robotics and embodied AI.
@@ 8 Reproducibility Statement
To ensure the reproducibility of our results, we have made our datasets, code, and experimental details available. All datasets used in this study are publicly available; we provide detailed descriptions of these datasets and all data preprocessing steps in Sec. 5.2 and Appendix A.2.1. The source code used for our experiments is included in the supplementary material. Upon acceptance, we will release the complete, documented source code under a permissive open-source license to facilitate the reproduction of all presented results. Key hyperparameters, model architectures, and training configurations are also detailed in Appendix A.2.1.
@@ 9 Use of LLM
Appendix A Appendix
A.1 Reward Design
A.1.1 Math reward design
We consider math QA with horizon $K$ and optional tool calls. Let $h_{k}$ be the dialogue/tool history at turn $k$. We adopt Math-Verifier [^2] Math-Verify (Hugging Face), GitHub: huggingface/Math-Verify. We use it as a parsing/normalization front-end and then apply a numeric comparator. as the checker front-end.
@@ A.1.1 Math reward design
$\textsc{NumEq}_{\delta}(a,b)\;=\;\mathbf{1}\!\left\{|a-b|\leq\delta\ \ \text{or}\ \ \frac{|a-b|}{\max(1,|b|)}\leq\delta\right\},\quad\delta{=}10^{-6}\ \text{(default)}.$
@@
Sparse pass at termination via numerical equality:
$r^{\mathrm{team}}_{k}\;=\;\mathbf{1}\{k{=}T,\ \textsc{CheckFinal}_{\textsc{MathVerifier}{+}\textsc{NumEq}}(h_{k}){=}\textsf{pass}\}\in\{0,1\},\qquad\lambda_{\mathrm{math}}=0.70.$
@@
Each agent $i$ uses a convex combination of verifiable sub-scores $s_{\ell,k}^{i}\in[0,1]$:
$r^{i,\mathrm{loc}}_{k}\;=\;\sum_{\ell\in\{\mathrm{fmt},\mathrm{tool},\mathrm{step}\}}c^{i}_{\ell}\,s^{i}_{\ell,k},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
@@
$s^{\mathrm{Reasoner}}_{\mathrm{fmt},k}=\mathbf{1}\{\text{required output schema matched at }k\},$
$s^{\mathrm{Reasoner}}_{\mathrm{step},k}=\begin{cases}\textsc{NumEq}_{\delta}\!\big(\hat{y}_{k},\,y^{\star}\big),&\text{if }\textsc{MathVerifier}\text{ extracts a numeric }\hat{y}_{k}\text{ at }k,\\[2.0pt] 0,&\text{otherwise},\end{cases}$
$m^{\mathrm{Reasoner}}_{k}=\mathbf{1}\{y^{\star}\ \text{available (}\textsc{MathVerifier}\text{)}\}.$
Tool (Python/calculator) local design.

### v4

@@ 4.2 MAS Training System
Mainstream RL post-training frameworks for LLMs, e.g., TRL (vonwerra2020trl), VERL (sheng2024hybridflow), AReaL (fu2025areal), and OpenRLHF (hu2024openrlhf) primarily support single-agent RL training, which typically involves: a single agent-environment interaction pattern, a single policy operating on a single data buffer, and a single LLM resource pool. This makes it difficult to (i) train multiple models in on-policy RL, (ii) maintain clean on-policy training data, and (iii) support diverse MAS workflow.
We introduce a novel MAS training system to overcome these challenges and enable AT-GRPO . By allocating an independent resource pool to each model, our system is designed to support the concurrent on-policy training of multiple policies. The system, depicted in Fig. 4, consists of the following components:
@@
Each policy is managed within an independent resource pool. Following HybridFlow-style  (sheng2025hybridflow), each pool comprises two workers: a RolloutWorker for inference and an UpdateWorker for optimization. During the rollout phase, all policies interact collectively according to the Alg. 1 and MAS workflow; Once collected, each trajectory is routed to the corresponding UpdateWorker, maintaining an on-policy learning regime for every policy.
@@ 5 Experiments
1. Experimental Setup. We train and evaluate Qwen3 models at 1.7B and 8B in the no-thinking mode (Qwen3TR). All runs use a single node with $8\times$ H100 GPUs. The rollout sample size is K=4K{=}4 and the turn horizon is $T=4$ for both multi-agent (MA) and single-agent (SA) settings. The reward-mixing coefficient is $\alpha{=}1$ without further tuning. Full training details appear in Appendix C.1.
@@
We evaluate five variants (all initialized from the same base model): (a) Single Agent (prompt-only): one frozen LLM solves the task end-to-end; (b) Single Agent + GRPO: as (a) but trained with GRPO (shao2024deepseekmath); (c) MAS (prompt-only): role-specialized prompting over a frozen, role-sharing backbone; (d) MAS + RL (role-sharing policy): all roles share one policy and pooled trajectories update it jointly; (e) MAS + RL (role-specialized policies): samples are routed by role and each policy is optimized independently (no parameter sharing).
3. Task Setups and Baselines.
To ensure a fair comparison, we align all environmental observations and reward signals across both MA and SA settings. While both paradigms utilize the same role-specific reward functions, the sole distinction is that the MA framework involves multiple agents capable of discussion. Detailed prompt templates and reward specifications are provided in Appendix B and C.2.
Single Agent Variants. We also evaluated a multi-turn SA variant for both Code and Math (see Appendix F), where a single agent repeatedly revises its own output until self-consistency. This setup is inherently less natural: the agent receives no additional environmental signal or cross-agent feedback, and the interaction pattern deviates from the QA-style pretraining regime of LLMs. Reflecting this mismatch, the multi-turn SA variant brought no empirical improvement and sometimes slightly degraded performance relative to the standard single-turn SA baseline.
Code. The environment observation is restricted to the problem description. The MA setting employs a dual-role debating mechanism: a Tester generates unit tests and a Coder generates the code. They refine their outputs iteratively until alignment is reached or the maximum turn limit is met. The natural SA baseline employs a Coder to generate the solution directly, as no other environmental feedback and other agent’s output is available.
Math. The environment observation consists is restricted to the problem. MA uses Dual-role debating MAS: a Tool-User (utilizing code interpreters) and a Reasoner (performing direct reasoning) until alignment or the maximum turn limit is met. The SA setting uses a Reasoner to derive the answer directly, as no other environmental feedback and other agents’ response.
Planning and Game. The environment observations are game states of the current turn. The MA setting employs a collaboration mechanism: a Tool-User (executing the tools) and an Executor (verifying tool outputs and executing actions). The SA employs an Executor to perform actions. Both settings share identical termination conditions based on goal satisfaction or the turn budget.
@@
Sudoku and Sokoban. We evaluate our method on gaming tasks: a $4{\times}4$ Sudoku and a $6{\times}6$ Sokoban. We use instances with an automatic checker, following the symbolic task setup of SymBench (Chen2025CodeSteer). To ensure a fair evaluation, we generate distinct training and validation sets using different random seeds and verify there is no overlap.
Plan-Path. We use a $10{\times}10$ grid-based Plan-Path environment. This follows the checker-backed symbolic task setup in CodeSteer’s SymBench (Chen2025CodeSteer). To separate training and validation, we generate the two splits with distinct random seeds and verify no duplication.
Code Generation. For training, we adopt size-specific corpora: the 1.7B Qwen model is trained on the APPS training split (introductory-difficulty subset) (Hendrycks2021APPS), while the 8B model is trained on CodeContests (DeepMind2024CodeContests). For model-generated code, we use the dataset’s golden unit tests to score correctness; for model-generated UT, we use the dataset’s golden reference solutions to compute the reward. For evaluation, we use three widely adopted coding benchmarks spanning interview-style and contest-style settings: APPS (Hendrycks2021APPS), LiveCodeBench-v6 (white2024livebench), and CodeContests (DeepMind2024CodeContests).
Mathematical Reasoning. We train on the Polaris-Dataset-53K (Polaris2025) and evaluate on several standard mathematical reasoning benchmarks. For validation, we use AIME24/AIME25 (AIME2024; AIME2025) and OlympiadBench (He2024OlympiadBench). All math tasks use verifier-checked numeric scoring.
@@ 5.2 Results and Analysis
MAS + AT-GRPO elevates the success rate from a 14–47% range for the single-agent baseline to 96.0–99.5%. By analyzing the dialogue records between agents, we find this dramatic improvement stems from an emergent collaboration: the tool agent learns to generate correct algorithms (e.g., BFS, $A^{\star}$ search), while the plan agent provides crucial oversight, interpreting execution outcomes and delivering the corrective final action list. On-policy RL training within the MAS enhances inter-agent coordination. Conversely, training agents in isolation results in only limited improvement, as detailed in our ablation study (Sec. 5.4, Tab. 4). Furthermore, on the coding and math benchmarks, our approach yields consistent gains, with absolute gains over the baseline ranging from +2.35 (CodeContests) to +16.30 (APPS) in coding, and from +1.80 (OlympaidBench) to +38.70 (AIME24) in math. We hypothesize two reasons: (1) Base models like Qwen3 have already been extensively trained for these common domains, as noted in their official reports (Qwen3TR), potentially leading to performance saturation. (2) The diverse nature of problems within these domains presents a greater challenge for improvement via RL training.
@@
Figure 5: (a) The system aggregates outputs from an ensemble of NN Reasoners and $M$ Tool-Users into a Judge. The total agent count scales as $M+N+1$, allowing for flexible resource allocation. (b) Evaluation on AIME24 (using Qwen3-8B).
Limitations of MAS-GRPO.
Empirical results in Tab. 2 and 2 indicate that directly applying GRPO to MAS often results in performance degradation. Notably, Qwen3-8B exhibits suboptimal results CodeContests (17.60 $\to$ 10.30) and OlympiadBench (56.50 $\to$ 53.20). We attribute this to the violation of the identical-state assumption: as multi-turn interaction histories diverge, the group-averaged baseline incorrectly aggregates heterogeneous states. This structural misalignment biases advantage estimation and destabilizes optimization.
Scalability Analysis with Collaborative Agents. To investigate scalability, we deploy a modular MAS architecture comprising Reasoners, Tool-Users, and a Judge (Fig. 5(a)). By varying the number of concurrent Reasoners ($N$) and Tool-Users ($M$), we scale the total agent count ($M+N+1$) to expand the exploration space. Fig. 5(b) demonstrates the scalability of MAS+AT-GRPO. While the baseline MAS+GRPO fails to scale effectively—saturating at 34.1% accuracy in the 7-agent regime—our method successfully leverages the increased ensemble size, achieving a continuous performance gain from 18.2% to 47.7%. This confirms that MAS+AT-GRPO can effectively scale across multiple agents without hitting the coordination bottlenecks observed in baselines. For a detailed analysis of computational efficiency and complexity, please refer to Appendix G.
Table 3: Comparison with existing MARL frameworks. We report Accuracy (%) for math/logic tasks and Pass@1 (%) for code tasks.
Math: Accuracy (%)
Backbone Method Config Acc.
Phi-3-mini Vanilla Baseline Zero-shot 65.0
MAPORL Trained 81.0
Ours (MAS) Untrained 84.4
Ours (MAS+AT-GRPO) Trained 88.7
Qwen2.5-Coder-3B-Instruct Vanilla Baseline Zero-shot 76.8
MARFT Trained 78.7
Ours (MAS) Untrained 84.4
Ours (MAS+AT-GRPO) Trained 87.1
Code: Pass@1 (%)
Backbone Method CodeContests LiveCodeBench
Qwen-2.5-7B-Instruct Vanilla Baseline 22.8 26.9
CURE 25.9 31.2
Ours (MAS) 30.3 30.4
Ours (+AT-GRPO) 34.2 35.3
Results cited from original papers.
5.3 Comparison with other MARL Frameworks
To assess the efficacy of our framework, we conducted ablation studies against representative baselines: As summarized in Tab. 3. we benchmark against three representative baselines: MAPORL (park2025maporl2), MARFT (Liao2025MARFT), and CURE wang2025cure. For fair comparison, we utilize identical base models and dataset splits. Our analysis highlights the advantages of two distinct MAS features: heterogeneous agent roles and multi-turn iterative interaction, as summarized in Tab. 3.
Comparison with MAPORL. We compare our approach with MAPORL using the Phi-3-mini-128k (3.4B) on gsm8k dataset cobbe2021training. While MAPORL relies on a debating mechanism among homogeneous agents, our framework realizes role heterogeneity—synergizing a Reasoning Agent with a Tool-use Agent for verification. This structural specialization proves superior: our untrained MAS achieves 84.4%, outperforming the trained MAPORL (81.0%). With AT-GRPO training, our performance further improves to 88.7%.
Comparison with MARFT and CURE. This comparison highlights the critical efficacy of iterative alignment over single-turn workflows. In math reasoning ( Qwen2.5-Coder-3B-Instruct), while MARFT relies on single-turn preference optimization, our framework leverages multi-turn interactions to facilitate active error correction and ambiguity resolution. Consequently, our inference-only MAS (84.4%) significantly outperforms the trained MARFT (78.7%), confirming that an extended reasoning horizon contributes more to robustness than single-step alignment; AT-GRPO training further amplifies this to 87.1%. A similar structural advantage is evident against CURE in code generation (Tab. 3). While CURE generates code and unit tests in a single turn, without utilizing them for self-correction. Our framework establishes a self-refinement cycle. This enables iterative debugging using generated tests, boosting CodeContests accuracy from 22.8% (vanilla) to 30.3% (untrained), surpassing the CURE baseline (25.9%), and ultimately reaching 34.2% with training.
@@ 5.4 Ablation Study
To further investigate the contributions of our core training components, We also conducted an ablation study with results summarized in Tab. 4 and Fig. 6. Our analysis yields several observations.
@@
Figure 6: (a) Evolution of standardized rewards for the Tool and Plan agents in the role-specific MAS on Plan-Path with Qwen3 1.7B. Shaded bands show variability across runs. (b) Evolution of the average turns required to solve tasks.
@@
6 Conclusion
In this paper, we proposed AT-GRPO, an agent- and turn-wise grouped reinforcement learning algorithm tailored for on-policy training in MAS. To support this, we introduced a novel training system capable of managing diverse MAS workflows and performing on-policy updates for multiple policies. Our extensive experiments demonstrate that our method delivers consistent gains across diverse domains. On planning tasks, it overcomes a key bottleneck of single-agent RL by boosting accuracy from a 14–47% baseline to 96.0–99.5%. Furthermore, it improves reasoning performance with average gains of 3.87–7.62% on coding and 9.0–17.93% on math tasks. Our analysis reveals that RL training in MAS context reinforces role-specific specialization, with the choice between a shared or specialized policy contingent on the task’s characteristics.
@@ 8 Reproducibility Statement
To ensure the reproducibility of our results, we have made our datasets, code, and experimental details available. All datasets used in this study are publicly available; we provide detailed descriptions of these datasets and all data preprocessing steps in Sec. 5.2 and Appendix C.1. The source code used for our experiments is included in the supplementary material. Upon acceptance, we will release the complete, documented source code under a permissive open-source license to facilitate the reproduction of all presented results. Key hyperparameters, model architectures, and training configurations are also detailed in Appendix C.1.
@@ 9 Use of LLM
Appendix A Extended Comparison of RL Training for LLM-based MAS
To complement the discussion on RL training for LLM-based multi-agent systems in Sec. 2, we provide an extended, axis-by-axis comparison of representative frameworks in Table 5. We contrast MAPoRL, MARFT, MARTI, CURE, and our method along several key design dimensions: (i) whether agents share a single policy or use role-specific policies; (ii) whether the interaction pattern is sequential, parallel, or a hybrid of both; (iii) whether the framework supports multi-turn interaction; (iv) whether agent roles are heterogeneous; (v) the number of evaluation task domains; and (vi) whether the underlying RL algorithm is designed as a general-purpose MAS training scheme rather than being tightly coupled to a single task.
Table 5: Comparison of RL-based LLM multi-agent training frameworks.
Method Policy sharing Execution pattern Multi-turn Role heterogeneity $\geq 2$ domains verification Generic MAS RL algo.
MAPoRL park2025maporl R P
MARFT Liao2025MARFT R S
MARTI zhang2025marti R S+P
CURE wang2025cure S P
Ours (StrongerMAS) S+R S+P
Note. For MARFT [Liao2025MARFT], we report the characteristics of version v3 of the framework, corresponding to the preprint available prior to the completion of this work.
As summarized in Tab. 5, StrongerMAS is the only RL-based LLM multi-agent training framework that simultaneously supports both shared and role-specific policies, hybrid sequential–parallel execution, multi-turn interaction, and heterogeneous agent roles, while being validated on multiple task domains and implemented as a general-purpose MAS RL algorithm. In contrast, MAPoRL [park2025maporl] and CURE [wang2025cure] restrict policies to a purely role-shared setting and are tailored to specific tasks, MARFT Liao2025MARFT restricts agents to single-turn sequential interactions, and MARTI zhang2025marti merely introduces basic single-agent RL algorithms (e.g., GRPO) to the MAS setting. . This combination of flexible workflow expressivity (S+R, S+P, heterogeneous roles) and broad, cross-domain evaluation makes StrongerMAS a more faithful and scalable abstraction for training cooperative LLM-based MAS.
Appendix B Reward Design
B.1 Math reward design
We consider math QA with horizon $T$ and optional tool calls. Let $h_{t}$ be the dialogue/tool history at turn $t$. We adopt Math-Verifier [^2] Math-Verify (Hugging Face), GitHub: huggingface/Math-Verify. We use it as a parsing/normalization front-end and then apply a numeric comparator. as the checker front-end.
@@ B.1 Math reward design
$\textsc{NumEq}_{\delta}(a,b)\;=\;\mathbf{1}\!\left\{|a-b|\leq\delta\ \ \text{or}\ \ \frac{|a-b|}{\max(1,|b|)}\leq\delta\right\},\quad\delta{=}10^{-6}\,.$
@@
Sparse pass at termination via numerical equality. We broadcast the same scalar reward to all turns:
$r^{\mathrm{team}}_{t}\;=\;\mathbf{1}\{\textsc{CheckFinal}_{\textsc{MathVerifier}{+}\textsc{NumEq}}(h){=}\textsf{pass}\}\in\{0,1\},\qquad\forall t.$
@@
Each agent $i$ at turn $t$ uses a masked convex combination of verifiable sub-scores $s^{i}_{\ell,t}\in[0,1]$:
$r^{\mathrm{loc}}_{t,i}\;=\;m_{t,i}\sum_{\ell\in\{\mathrm{fmt},\mathrm{tool},\mathrm{step}\}}c^{i}_{\ell}\,s^{i}_{\ell,t},\qquad\sum_{\ell}c^{i}_{\ell}=1,$
where $m_{t,i}\in\{0,1\}$ is a verifiability mask.
@@
$s^{\mathrm{Reasoner}}_{\mathrm{fmt},t}=\mathbf{1}\{\text{required output schema matched at turn }t\},$
$s^{\mathrm{Reasoner}}_{\mathrm{step},t}=\begin{cases}\textsc{NumEq}_{\delta}\!\big(\hat{y},\,y^{\star}\big),&\text{if }\textsc{MathVerifier}\text{ extracts a numeric }\hat{y},\\[2.0pt] 0,&\text{otherwise},\end{cases}$
and we do not use a tool-related step, so the corresponding score is implicitly $s^{\mathrm{Reasoner}}_{\mathrm{tool},t}\equiv 0$. The mask is
$m^{\mathrm{Reasoner}}_{t}=\mathbf{1}\{y^{\star}\ \text{available (}\textsc{MathVerifier}\text{) at turn }t\}.$
For the Reasoner, $r^{\mathrm{loc}}_{t,\mathrm{Reasoner}}$ is obtained by plugging these coefficients, scores and the mask into the generic form above.
B.2 Code reward design
We consider code synthesis with unit tests. Let $\mathcal{S}$ be the active test suite and
$p=\frac{1}{|\mathcal{S}|}\sum_{t\in\mathcal{S}}\mathbf{1}\{\textsc{Run}(t,\text{code})=\textsf{pass}\}\in[0,1].$
Team reward is dense; we again broadcast it over turns:
$r^{\mathrm{team}}_{t}\;=\;p,\qquad\forall t.$
Local rewards.
Each agent $i$ at turn $t$ uses a masked convex combination of verifiable sub-scores $s^{i}_{\ell,t}\in[0,1]$:
$r^{\mathrm{loc}}_{t,i}\;=\;m_{t,i}\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,t},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
Coder local reward.
We define the Coder’s local reward as a weighted combination of basic sanity checks and the fraction of golden tests passed by the generated code. Coefficients:
Let $\mathcal{T}^{\text{gold}}$ be the fixed set of golden unit tests for this problem. Component scores at turn $t$ are
[…]
$s^{\mathrm{Coder}}_{\mathrm{nr},t}=\frac{1}{|\mathcal{T}^{\text{gold}}|}\sum_{u\in\mathcal{T}^{\text{gold}}}\mathbf{1}\{\textsc{Run}(u,\text{code})=\textsf{pass}\},$
i.e., the fraction of golden tests passed by the current code code. We apply an availability mask
[…]
and define the Coder’s local reward as
$r^{\mathrm{loc}}_{t,\mathrm{Coder}}=m^{\mathrm{Coder}}_{t}\Big(c^{\mathrm{Coder}}_{\mathrm{build}}s^{\mathrm{Coder}}_{\mathrm{build},t}+c^{\mathrm{Coder}}_{\mathrm{run}}s^{\mathrm{Coder}}_{\mathrm{run},t}+c^{\mathrm{Coder}}_{\mathrm{nr}}s^{\mathrm{Coder}}_{\mathrm{nr},t}\Big).$
Tester local design.